OpenStudy (anonymous):

write (x+7)^2 in expanded fourm

OpenStudy (anonymous):

let's go step by step

OpenStudy (anonymous):

kay

OpenStudy (anonymous):

(x+7)^2 = (x+7)*(x+7) = x*x + x*7 + 7*x + 7*7 = x^2 + 7x + 7x + 49 = x^2 + 14x + 49

OpenStudy (anonymous):

understand?

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

in general: \((a+b)^2 = a^2 + 2ab + b^2\)

OpenStudy (anonymous):

got,ot

OpenStudy (anonymous):

and \((a-b)^2 = a^2 - 2ab + b^2\)

OpenStudy (anonymous):

*it